A positive charge of 1.000μ C is located in a uniform field of 8.00×104 N/C. A negative charge of -0.500μ C is brought near enough to the positive charge that the attractive force between the charges just equals the force on the positive charge due to the field. How close are the two charges?
Answer #1

Here ,

let the distance between the charges is d

as the force on the first charge due to second charge is equal to force on first charge due to electric field

the field produced by the second charge will be same as the electric field present

hence ,

k * q2/d^2 = 8 *10^4

9 *10^9 * 0.50 *10^-6/d^2 = 8 *10^4

d = 0.237 m

the distance between the charges is 0.237 m